# 3D Diffuser Actor: Multi-task 3D Robot Manipulation with Iterative Error Feedback

### v2

@@ III-A Denoising Diffusion Probabilistic Models
Observation space and action representation 3D Diffuser Actor is trained on demonstrations of successful trajectories in the form of $\{(o_{1},\mathbf{a}_{1}),(o_{2},\mathbf{a}_{2}),...\}$, accompanied with a task language description, similar to previous works [31, 69, 19, 80]. Each observation $o$ comprises a set of posed (multi-view) RGB-D images. Each action $\mathbf{a}$ describes an end-effector pose and is decomposed into 3D position, 3D orientation and a binary open/closed state: $\mathbf{a}=\{\mathbf{a}^{\mathrm{\small pos}}\in\mathbb{R}^{3},\mathbf{a}^{\mathrm{\small rot}}\in\mathbb{R}^{6},\mathbf{a}^{\mathrm{\small open}}\in\{0,1\}\}$. We represent rotations using the 6D rotation representation of [85] for all environments in all our experiments, to avoid the discontinuities of the quaternion representation.
@@
Keyposes 3D Diffuser Actor inherits the temporal abstraction of demonstrations into end-effector keyposes from previous works [31, 69, 19, 42]. Keyposes are important intermediate end-effector poses that summarize a demonstration and can be extracted using simple heuristics, such as a change in the open/close end-effector state or local extrema of velocity/acceleration. We describe the heuristics we use for each dataset in the Appendix (Section A-A). During inference, 3D Diffuser Actor can either predict and execute the full trajectory of actions up to the next keypose (including the keypose), or just predict the next keypose and use a sampling-based motion planner to reach it, similar to previous works [69, 23, 19]. In the rest of this section we use the term “trajectory" to refer to 3D Diffuser Actor’s output. Keypose prediction is a special case where the trajectory contains only one future pose.
Scene and language encoder We use a scene and language encoder similar to [19, 81]. We describe it here to make the paper self-contained. The input to our scene encoder is a set of posed RGB-D images. We first extract multi-scale visual tokens for each camera view using a pre-trained 2D feature extractor and a feature pyramid network [41]. Next, we associate every 2D feature grid location in the 2D feature maps with a depth value, by averaging the depth values of the image pixels that correspond to it. We use camera intrinsics and the pinhole camera equation to map a pixel location and depth value $(x,y,d)$ to a 3D location $(X,Y,Z)$, and “lift” the 2D feature tokens to 3D, to obtain a 3D feature cloud. The language encoder maps language task descriptions or instructions into language feature tokens. We use the pre-trained CLIP ResNet50 2D image encoder  [58] to encode each RGB image into a 2D feature map and the pre-trained CLIP language encoder to featurize the language task instruction.
3D Relative Position Denoising Transformer 3D Diffuser Actor iteratively denoises an estimate of the end-effector’s future trajectory. Every trajectory step token is a 10-dimensional action, as defined earlier. We project each such action into a high-dimensional representation through an MLP. Each action token comes with the positional embeddings of its corresponding $\mathbf{a}^{\mathrm{\small pos}}$. This enables relative cross-attention with the scene and other physical entities. We incorporate proprioceptive information as a short history of predicted end-effector keyposes. These are represented with learnable feature vectors and their 3D positions are used to compute positional embeddings.
We contextualize all tokens, namely visual $o$, language $l$, proprioception $c$ and action tokens (current estimate) $\mathbf{a}_{t}^{pos},\mathbf{a}_{t}^{rot}$ using 3D relative position attention layers. Inspired by recent work in visual correspondence [39, 21] and 3D manipulation [19, 81], we use rotary positional embeddings [74] with the property that the dot product of two positionally encoded features $\mathbf{x}_{i},\mathbf{x}_{j}$ is:
$\mathbf{PE}(\mathbf{p}_{i},\mathbf{x}_{i})^{T}\mathbf{PE}(\mathbf{p}_{j},\mathbf{x}_{j})=\mathbf{x}_{i}^{T}\mathbf{M}(\mathbf{p}_{j}-\mathbf{p}_{i})\mathbf{x}_{j}$ (3)
where $\mathbf{M}$ is a matrix function which depends only on the relative positions of points $\mathbf{p}_{i}$ and $\mathbf{p}_{j}$ and thus is translation-invariant. The updated action feature tokens are fed to MLPs to predict the position error $\epsilon_{\theta}^{\mathrm{\small pos}}(o,l,c,\mathbf{a}_{t}^{\mathrm{\small pos}},\mathbf{a}_{t}^{\mathrm{\small rot}},t)$, rotation error $\epsilon_{\theta}^{\mathrm{\small rot}}(o,l,c,\mathbf{a}_{t}^{\mathrm{\small pos}},\mathbf{a}_{t}^{\mathrm{\small rot}},t)$ and whether the end-effector should be open or closed $f_{\theta}^{\mathrm{\small open}}(o,l,c,\mathbf{a}_{t}^{\mathrm{\small pos}},\mathbf{a}_{t}^{\mathrm{\small rot}},t)$.
We use a modified version of Equation 2 to update the current estimate of each element of the end-effector’s trajectory:
$\mathbf{a}^{\mathrm{\small pos}}_{t-1}=$ $\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{a}^{\mathrm{\small pos}}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}^{\mathrm{\small pos}}(o,l,c,\mathbf{a}_{t}^{\mathrm{\small pos}},\mathbf{a}_{t}^{\mathrm{\small rot}},t)\right)$
$+\frac{1-\bar{\alpha}_{t+1}}{1-\bar{\alpha}_{t}}\beta_{t}\mathbf{z}^{\small pos}$ (4)
$\mathbf{a}^{\mathrm{\small rot}}_{t-1}=$ $\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{a}^{\mathrm{\small rot}}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}^{\mathrm{\small rot}}(o,l,c,\mathbf{a}_{t}^{\mathrm{\small rot}},\mathbf{a}_{t}^{\mathrm{\small rot}},t)\right)$
$+\frac{1-\bar{\alpha}_{t+1}}{1-\bar{\alpha}_{t}}\beta_{t}\mathbf{z}^{\small rot}$ (5)
where $\mathbf{z}^{\small pos},\mathbf{z}^{\small rot}\sim\mathcal{N}(\mathbf{0},\mathbf{1})$ variables of appropriate dimension.
We use the following two noise schedulers:
  1. 1.
    a scaled-linear noise scheduler $\beta_{t}=(\beta_{\max}-\beta_{\min})t+\beta_{\min}$, where $\beta_{\max},\beta_{\min}$ are hyperparameters, set to $0.02$ and $0.0001$ in our experiments,
  2. 2.
    a squared cosine noise scheduler $\beta_{t}=\frac{1-\cos\big{(}\frac{(t+1)/T+0.008}{1.008}*\frac{\pi}{2}\big{)}^{2}}{\cos\big{(}\frac{t/T+0.008}{1.008}*\frac{\pi}{2}\big{)}^{2}}$.
We found using a scaled-linear noise scheduler for denoising end-effector’s 3D positions and a squared cosine noise scheduler for denoising the end-effector’s 3D orientations to converge much faster than using squared cosine noise for both.
Training 3D Diffuser Actor is trained on a dataset of kinesthetic demonstrations, which consists of tuples of RGB-D observations, proprioception information of robot’s end-effector pose $c$, action trajectories $\mathbf{a}=[\mathbf{a}^{pos},\mathbf{a}^{rot},\mathbf{a}^{open}]$ and corresponding language goals $\mathcal{D}=\{(o_{1},c_{1},\mathbf{a}_{1},l_{1}),(o_{2},c_{2},\mathbf{a}_{2},l_{2}),...\}$. During training, we randomly sample a diffusion step $t$ and add noise $\epsilon=(\epsilon^{\mathrm{\small pos}}_{t},\epsilon^{\mathrm{\small rot}}_{t})$ to the ground-truth target poses (position and rotation). We supervise 3D Diffuser Actor using a denoising objective: conditioning on the tuple of $(o,l,c,t)$, our model learns to reconstruct the clean actions by predicting the pose transformation with respect to the current estimate. We use the $L1$ loss for reconstructing the 3D position and 3D rotation error. We use binary cross-entropy loss to supervise end-effector opening $f_{\theta}^{\mathrm{\small open}}$. Our objective reformulates Equation 1 into:
$\mathcal{L}_{\theta}=$ $\frac{1}{|\mathcal{D}|}\sum_{i=1}^{|\mathcal{D}|}\mathrm{BCE}(f_{\theta}^{\mathrm{\small open}}(o_{i},l_{i},c_{i},\mathbf{a}_{t,i}^{\mathrm{\small pos}},\mathbf{a}_{t,i}^{\mathrm{\small rot}},t),\mathbf{a}_{i}^{\mathrm{\small open}})$
$+w_{1}\cdot\|(\epsilon_{\theta}^{\mathrm{\small pos}}(o_{i},l_{i},c_{i},\mathbf{a}_{t,i}^{\mathrm{\small pos}},\mathbf{a}_{t,i}^{\mathrm{\small rot}},t)-\epsilon_{t}^{\mathrm{\small pos}}\|$
$+w_{2}\cdot\|(\epsilon_{\theta}^{\mathrm{\small rot}}(o_{i},l_{i},c_{i},\mathbf{a}_{t,i}^{\mathrm{\small pos}},\mathbf{a}_{t,i}^{\mathrm{\small rot}},t)-\epsilon^{\mathrm{\small rot}}_{t}\|,$ (6)
where $w_{1},w_{2}$ are hyperparameters estimated using cross-validation.
Implementation details We lift 2D image features to 3D by calculating xyz-coordinates of each image pixel, using the sensed depth and camera parameters. We augment RGB-D observations with random rescaling and cropping. Nearest neighbor interpolation is used for rescaling RGB-D observations. To reduce the memory footprint in our 3D Relative Transformer, we use Farthest Point Sampling to sample a subset of the points in the input 3D feature cloud. We use FiLM [55] to inject conditional input, including the diffusion step and proprioception history, to every attention layer in the model. We include a detailed architecture diagram of our model and a table of hyper-parameters used in our experiments in the Appendix (Section A-E and Section A-F).
IV Experiments
We test 3D Diffuser Actor in learning multi-task manipulation policies on RLBench [32] and CALVIN [49], two established learning from demonstrations benchmarks. We compare against the current state-of-the-art on single-view and multi-view setups, as well as against ablative versions of our model that do not consider 3D scene representations or relative attention. We also test 3D Diffuser Actor’s ability to learn multi-task policies from real-world demonstrations.
IV-A Evaluation on RLBench
Setup RLBench is built atop the CoppelaSim [61] simulator, where a Franka Panda Robot is used to manipulate the scene. Our 3D Diffuser Actor is trained to predict the next end-effector keypose and we employ the low-level motion planner BiRRT [37], native to RLBench, to reach the predicted pose, following previous works [69, 19]. Our model does not perform collision checking for our experiments. We train and evaluate 3D Diffuser Actor on two experimental setups of multi-task manipulation:
  1. 1.
    PerAct setup introduced in [69]: This uses a suite of 18 manipulation tasks, each task has 2-60 variations, which concern scene variability across object poses, appearance and semantics. The tasks are specified by language descriptions. There are four cameras available (front, wrist, left shoulder, right shoulder). There are 100 training demonstrations available per task, evenly split across task variations and 25 unseen test episodes for each task. Due to the randomness of the sampling-based motion planner, we test our model across 5 random seeds. During evaluation, models are allowed to predict a maximum of 25 keyposes, unless they receive earlier task-completion indicators from the simulation environment.
  2. 2.
    GNFactor setup introduced in [83]: This uses a suite of 10 manipulation tasks (a subset of PerAct’s task set). Only one RGB-D camera view is available (front camera). There are 20 training demonstrations per task, evenly split across variations. The evaluation setup mirrors that of PerAct, with the exception that we test the final checkpoint using 3, rather than 5, random seeds on the test set.
Baselines We consider the following baselines:
  1. 1.
    C2F-ARM-BC [33], a 3D policy that iteratively voxelizes RGB-D images and predicts actions in a coarse-to-fine manner. Q-values are estimated within each voxel and the translation action is determined by the centroid of the voxel with the maximal Q-values.
  2. 2.
    PerAct [69], a 3D policy that voxelizes the workspace and detects the next voxel action through global self-attention.
  3. 3.
    Hiveformer [23], a 3D policy that enables attention between features of different history time steps.
  4. 4.
    PolarNet [8], a 3D policy that computes dense point representations for the robot workspace using a PointNext backbone [57].
  5. 5.
    RVT [22], a 3D policy that deploys a multi-view transformer to predict actions and fuses those across views by back-projecting to 3D.
  6. 6.
    Act3D  [19], a 3D policy that featurizes the robot’s 3D workspace using coarse-to-fine sampling and featurization.
  7. 7.
    GNFactor [83], a 3D policy that co-optimizes a neural field for reconstructing the 3D voxels of the input scene and a PerAct module for predicting actions based on voxel representations.
We report results for RVT, PolarNet and GNFactor based on their respective papers. Results for CF2-ARM-BC and PerAct are presented as reported in RVT [22]. Results for Hiveformer are copied from PolarNet [8].
We observed that Act3D [19] does not follow the same setup as PerAct on RLBench. Specifically, Act3D uses different 1) 3D object models, 2) success conditions, 3) training/test episodes and 4) maximum numbers of keyposes during evaluation. For fair comparison, we retrain and test Act3D on the same setup.
We also compare to the following ablative versions of our model:
  1. 1.
    2D Diffuser Actor, our implementation of [11]. We remove the 3D scene encoding from 3D Diffuser Actor and instead use per-image 2D representations by average-pooling features within each view. We add learnable embeddings to distinguish between different views. We use standard attention layers for joint encoding the action estimate and 2D image features.
  2. 2.
    3D Diffuser Actor w/o Rel. Attn., an ablative version of our model that uses standard (non-relative) attentions to featurize the current rotation and translation estimate with the 3D scene feature cloud. This version of our model is not translation-equivariant.
Evaluation metrics Following previous work [19, 69], we evaluate policies by task completion success rate, which is the proportion of execution trajectories that achieve the goal conditions specified in the language instructions.

### v3

@@ 3 Method
3D Diffuser Actor is trained to imitate demonstration trajectories of the form of $\{(\mathbf{o}_{1},\mathbf{a}_{1}),(\mathbf{o}_{2},\mathbf{a}_{2}),...\}$, accompanied with a task language instruction $l$, similar to previous works [james2022q, shridhar2023perceiver, gervet2023act3d, wu2023unleashing], where $\mathbf{o}_{t}$ stands for the visual observation and $\mathbf{a}_{t}$ stands for robot action at timestep $t$. Each observation $\mathbf{o}_{t}$ is one or more posed RGB-D images. Each action $\mathbf{a}_{t}$ is an end-effector pose and is decomposed into 3D location, rotation and binary (open/close) state: $\mathbf{a}_{t}=\{\mathbf{a}_{t}^{\mathrm{loc}}\in\mathbb{R}^{3},\mathbf{a}_{t}^{\mathrm{rot}}\in\mathbb{R}^{6},\mathbf{a}_{t}^{\mathrm{open}}\in\{0,1\}\}$. We represent rotations using the 6D rotation representation of [zhou2020continuity] for all environments in all our experiments, to avoid the discontinuities of the quaternion representation. We will use the notation $\bm{\tau}_{t}=(\mathbf{a}_{t:t+T}^{\mathrm{loc}},\mathbf{a}_{t:t+T}^{\mathrm{rot}})$ to denote the trajectory of 3D locations and rotations at timestep $t$, of temporal horizon $T$. Our model, at each timestep $t$ predicts a trajectory $\bm{\tau}_{t}$ and binary states $\mathbf{a}_{t:t+T}^{\mathrm{open}}$.
The architecture of 3D Diffuser Actor is shown in Figure 1. It is a conditional diffusion probabilistic model [sohldickstein2015deep, ho2020denoising] of trajectories given the visual scene and a language instruction; it predicts a whole trajectory $\bm{\tau}$ at once, non autoregressively, through iterative denoising, by inverting a process that gradually adds noise to a sample $\bm{\tau}^{0}$. The diffusion process is associated with a variance schedule $\{\beta^{i}\in(0,1)\}_{i=1}^{N}$, which defines how much noise is added at each diffusion step. The noisy version of sample $\bm{\tau}^{0}$ at diffusion step $i$ can then be written as $\bm{\tau}^{i}=\sqrt{\bar{\alpha}^{i}}\bm{\tau}^{0}+\sqrt{1-\bar{\alpha}^{i}}\epsilon$, where $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{1})$ is a sample from a Gaussian distribution (with the same dimensionality as $\bm{\tau}^{0}$), $\alpha^{i}=1-\beta^{i}$, and $\bar{\alpha}^{i}=\prod_{j=1}^{i}\alpha^{j}$.
3D Relative Denoising Transformer 3D Diffuser Actor models a learned gradient of the denoising process with a 3D relative transformer $\hat{\epsilon}=\epsilon_{\theta}(\bm{\tau}_{t}^{i};i,\mathbf{o}_{t},l,c_{t})$ that takes as input the noisy trajectory $\bm{\tau}_{t}^{i}$ at timestep $t$, diffusion step $i$, and conditioning information from the language instruction $l$, the visual observation $\mathbf{o}_{t}$ and proprioception $c_{t}$ of timestep $t$, to predict the noise component $\hat{\epsilon}$. At each timestep $t$ and diffusion step $i$, we convert visual observations $\mathbf{o}_{t}$, proprioception $c_{t}$ and noised trajectory estimate $\bm{\tau}_{t}^{i}$ to a set of 3D tokens. Each 3D token is represented by a latent embedding and a 3D position. Our model fuses all 3D tokens using relative 3D attentions, and additionally fuses information from the language instruction using normal attentions, since it is meaningless to define 3D coordinates for language tokens. Next, we describe how we featurize each piece of input (when not ambiguous, we omit the subscript $t$ for clarity).
3D tokenization At each diffusion step $i$, we represent the noisy estimate $\bm{\tau}^{i}$ of the clean trajectory $\bm{\tau}^{0}$ as a sequence of 3D trajectory tokens, by mapping each noisy pose $\mathbf{a}^{i}$ of $\bm{\tau}^{i}$, to a latent embedding vector with a MLP, and a 3D position which comes from the noised 3D translation component $\mathbf{a}^{loc,i}$. We featurize each image view using a 2D feature extractor and obtain a corresponding 2D feature map $F\in\mathbb{R}^{H\times W\times c}$, where $c$ is the number of feature channels and $H,W$ the spatial dimensions, using a pre-trained CLIP ResNet50 2D image encoder  [radford2021learning]. Given the corresponding depth map from that view, we compute the 3D location $(X,Y,Z)$ for each of the $H\times W$ feature patches by averaging the depth value inside the patch extent. We map the pixel coordinate and depth value of a patch to a corresponding 3D coordinate using camera intrinsics and extrinsics and the pinhole camera model. This results in a 3D scene token set of cardinality $H\times W$. Each scene token is represented by the corresponding patch feature vector $F_{x,y}$, the feature vector at the corresponding patch coordinate $(x,y)$, and a 3D position. If more than one views are available, we aggregate 3D scene tokens from each, to obtain the final 3D scene token set. The proprioception $c$ is also a 3D scene token, with a learnable latent representation and the 3D positional embedding corresponding to the end-effector’s current 3D location. Lastly, we map the language task instruction to language tokens using a pre-trained CLIP language encoder, following previous works [gervet2023act3d].
Our 3D Relative Denoising Transformer applies relative self-attentions among all 3D tokens, and cross-attentions to the language tokens. For the 3D self-attentions, we use the rotary positional embeddings [su2021roformer] to encode relative positional information in attention layers. The attention weight between query $q$ and key $k$ is written as: $e_{q,k}\propto\mathbf{x}_{q}^{T}\mathbf{M}(\mathbf{p}_{q}-\mathbf{p}_{k})\mathbf{x}_{k}$, where $\mathbf{x}_{q}$ / $\mathbf{x}_{k}$ and $\mathbf{p}_{q}$ / $\mathbf{p}_{k}$ denote the features and 3D positions of the query / key, and $\mathbf{M}$ is a matrix function which depends only on the relative positions of the query and key, inspired by recent work in visual correspondence [Li2021LepardLP, Gkanatsios2023AnalogyFormingTF] and 3D manipulation [gervet2023act3d, xian2023chaineddiffuser]. We feed the final trajectory tokens to MLPs to predict: (1) the noise $\epsilon^{loc}_{\theta}(\mathbf{o},l,c,\bm{\tau}^{i},i)$ and $\epsilon^{rot}_{\theta}(\mathbf{o},l,c,\bm{\tau}^{i},i)$ added to $\bm{\tau}^{0}$’s sequence of 3D translations and 3D rotations, respectively, and (2) the end-effector opening $f_{\theta}^{\mathrm{open}}(\mathbf{o},l,c,\bm{\tau}^{i},i)\in[0,1]^{T}$.
Training and inference During training, we randomly sample a time step $t$ and a diffusion step $i$ and add noise $\epsilon=(\epsilon^{\mathrm{loc}},\epsilon^{\mathrm{rot}})$ to a ground-truth trajectory $\bm{\tau}_{t}^{0}$. We use the $L1$ loss for reconstructing the sequence of 3D locations and 3D rotations. We use binary cross-entropy (BCE) loss to supervise the end-effector opening $f_{\theta}^{\mathrm{open}}$, we use the prediction from i=1 at inference time. Our objective reads:
$\mathcal{L}_{\theta}=w_{1}\|(\epsilon_{\theta}^{\mathrm{loc}}(\mathbf{o},l,c,\bm{\tau}^{i},i)-\epsilon^{\mathrm{loc}}\|+w_{2}\|(\epsilon_{\theta}^{\mathrm{rot}}(\mathbf{o},l,c,\bm{\tau}^{i},i)-\epsilon^{\mathrm{rot}}\|+\mathrm{BCE}(f_{\theta}^{\mathrm{open}}(\mathbf{o},l,c,\bm{\tau}^{i},i),\mathbf{a}_{1:T}^{\mathrm{open}}),$
where $w_{1},w_{2}$ are hyperparameters estimated using cross-validation. To draw a sample from the learned distribution $p_{\theta}(\bm{\tau}|\mathbf{o},l,c)$, we start by drawing a sample $\bm{\tau}_{N}\sim\mathcal{N}(\mathbf{0},\mathbf{1})$. Then, we progressively denoise the sample by iterated application of $\epsilon_{\theta}$ $N$ times according to a specified sampling schedule [DDPM, song2020denoising], which terminates with $\bm{\tau}_{0}$ sampled from $p_{\theta}(\bm{\tau})$: $\bm{\tau}^{i-1}=\frac{1}{\sqrt{\alpha^{i}}}\left(\bm{\tau}^{i}-\frac{\beta^{i}}{\sqrt{1-\bar{\alpha}_{i}}}\epsilon_{\theta}(\mathbf{o},l,c,\bm{\tau}^{i},i)\right)+\frac{1-\bar{\alpha}^{i+1}}{1-\bar{\alpha}^{i}}\beta^{i}\mathbf{z},$ where $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{1})$ is a random variable of appropriate dimension. Empirically, we found that using separate noise schedulers for $\mathbf{a}^{\mathrm{loc}}$ and $\mathbf{a}^{\mathrm{rot}}$, specifically, scaled-linear and square cosine scheduler, respectively, achieves better performance.
@@
Implementation details At training, we segment demonstration trajectories at detected end-effector keyposes, such as a change in the open/close end-effector state or local extrema of velocity/acceleration, following previous works [james2022q, shridhar2023perceiver, gervet2023act3d, liu2022instruction]. We then resample each trajectory segment to have the same length $T$. During inference, 3D Diffuser Actor can either predict and execute the full trajectory of actions up to the next keypose (including the keypose), or just predict the next keypose and use a sampling-based motion planner to reach it, similar to previous works [shridhar2023perceiver, guhur2023instruction, gervet2023act3d].
Due to space limits, please check Section B.3 for detailed model diagram, Section B.5 for our choice of hyper-parameters, Section B.6 for detailed formulation of denoising diffusion probabilistic models and Section B.7 for discussion of noise schedulers in the Appendix.
4 Experiments
We test 3D Diffuser Actor in multi-task manipulation on RLBench [james2020rlbench] and CALVIN [mees2022calvin], two established learning from demonstrations benchmarks, and in the real world.
4.1 Evaluation on RLBench
RLBench is built atop the CoppelaSim [rohmer2013v] simulator, where a Franka Panda Robot is used to manipulate the scene. On RLBench, our model and all baselines are trained to predict the next end-effector keypose as opposed to keypose trajectory; all methods employ a low-level motion planner BiRRT [kuffner2000rrt], native to RLBench, to reach the predicted robot keypose. We train and evaluate 3D Diffuser Actor on two setups based on the number of available cameras: 1. Multi-view setup, introduced in [shridhar2023perceiver], that uses a suite of 18 manipulation tasks, each with 2-60 variations, which concern scene variability across object poses, appearance and semantics. There are four RGB-D cameras available, front, wrist, left shoulder and right shoulder. The wrist camera moves during manipulation. 2. Single-view setup, introduced in [ze2023gnfactor], that uses a suite of 10 manipulation tasks. Only the front RGB-D camera view is available. We evaluate policies by task completion success rate, which is the proportion of execution trajectories that achieve the goal conditions specified in the language instructions [gervet2023act3d, shridhar2023perceiver].
Baselines All compared methods on RLBench are 3D policies that use depth and camera parameters during featurization of the RGB-D input. We compare against the following: C2F-ARM-BC [james2022coarse] and PerAct [shridhar2023perceiver] that voxelize the 3D workspace, Hiveformer [guhur2023instruction] that featurizes XYZ coordinates aligned with the 2D RGB frames, PolarNet [Chen2023PolarNet3P] that featurizes a scene 3D point cloud, GNFactor [ze2023gnfactor] that uses a single RGB-D view and is trained to complete the 3D feature volume, RVT [goyal2023rvt] and Act3D  [gervet2023act3d], that are the previous SOTA methods on RLBench. We report results for RVT, PolarNet and GNFactor based on their respective papers. Results for CF2-ARM-BC and PerAct are presented as reported in [goyal2023rvt]. Results for Hiveformer are copied from [Chen2023PolarNet3P]. We retrained Act3D on the multi-view setup using the publicly available code, because we found some differences on the train and test splits used in the original paper. We also trained Act3D on the single-view setup to use as additional baseline for the single-view setup, alongside GNFactor. 3D Diffuser Actor and all baselines are trained on the same set of keyposes extracted from expert demonstrations [james2022q].